OpenStudy (anonymous):

(2x - 3)(3x + 1) = 6x^2 + 2x - 9x - 3 = 6x^2 - 7x - 3

hero (hero):

(2x - 3)(3x + 1) = 3x(2x-3) + 1(2x-3) = 6x^2 - 9x + 2x - 3 = 6x^2 - 7x - 3

hero (hero):

That's how it is done, but otherwise, correct.